# A flexible Bayesian g-formula for causal survival analyses with time-dependent confounding

### v3

@@ 1 Introduction
Although the original form of the g-formula is nonparametric, its parametric implementation, based on the maximum likelihood estimates of its component models, has been more commonly used to ensure feasible computation. For instance, by employing generalized linear models, the parametric g-formula estimator has been implemented to adjust for a relatively large number of confounders with either non-censored (Westreich et al.,, 2012) or censored outcomes (Keil et al.,, 2014). However, the parametric g-formula estimator faces two potential challenges. First, for studies involving a moderate number of time-varying confounders, modeling the entire history of time-varying confounders can be a daunting task, especially when the time-varying confounders involve a mix of continuous and categorical variables (Achy-Brou et al.,, 2010). Second, the consistency of the parametric g-formula estimator requires the correct specification of all component models, whereas the need for modeling multi-dimensional confounders in each period is susceptible to model misspecification (Daniel et al.,, 2013). For example, overly parsimonious models can lead to the g-null paradox (Robins and Wasserman,, 1997) and introduce bias in the final causal effect estimates; see McGrath et al., (2022) for a further discussion. To this end, methods that improve the parametric g-formula with attention to more flexible model specifications are of increasing interest.
@@
Our second contribution is to formally integrate the Bayesian Additive Regression Trees (BART), a flexible machine learner, into the survival g-formula to improve upon the standard implementation of the survival parametric g-formula estimator in longitudinal observational studies. Specifically, we incorporate BART for component models in a g-formula to mitigate the bias arising from parametric model misspecification (Chipman et al.,, 2010; Sparapani et al.,, 2016; Tan and Roy,, 2019). We consider this approach because causal estimators based on BART have been empirically demonstrated to outperform competitors in drawing causal inference about the average treatment effect (Hill,, 2011), conditional average treatment effect (Henderson et al.,, 2018; Dorie et al.,, 2019; Hu et al.,, 2021; Hahn et al.,, 2020; Caron et al.,, 2022), as well as causal effects with an intermediate variable (Josefsson and Daniels,, 2021; Bargagli-Stoffi et al.,, 2022; Linero and Zhang,, 2022; Chen et al.,, 2024). This BART approach is flexible as it also provides full posterior samples for all unknown parameters, with built-in uncertainty quantification (Woody et al.,, 2021).
@@ 2.3 Assumptions and the g-formula—a brief review
Let $f(C_{t+1}=0|\overline{l}_{t},\overline{a}_{t}^{g},C_{t}=Y_{t}=0)$ denote the conditional probability of no censoring in period $t+1$, and then the joint density of treatment status and no censoring, $f(a_{t}^{g},C_{t+1}=0|\overline{l}_{t},\overline{a}_{t-1}^{g},C_{t}=Y_{t}=0)=f(a_{t}^{g}|\overline{h}_{t},\overline{a}_{t-1}^{g},C_{t}=Y_{t}=0)\times f(C_{t+1}=0|\overline{l}_{t},\overline{a}_{t}^{g},C_{t}=Y_{t}=0)>0$, whenever $f(\overline{L}_{t}=\overline{l}_{t},\overline{A}_{t}=\overline{a}_{t}^{g},C_{t}=Y_{t}=0)>0,\forall~{}t\in\mathscr{T}$.
@@ 3 A general class of survival g-formulas
Different specifications of the longitudinal balancing scores can potentially serve different purposes. For example, one can choose the longitudinal joint score or any bijective mapping of these scores as the longitudinal balancing score in (3). This approach could be preferable in some cases because $e_{s}$ is a scalar summary of the higher-dimensional $L_{s}$ in each period $s$, and it is substantially easier to model and then integrate over the one-dimensional longitudinal joint score than multi-dimensional confounders. For a dynamic treatment strategy, one could set $\overline{b}_{s}=\{\overline{H}_{s},\overline{e}_{s}\}$ and maintain a lower dimensionality relative to $\overline{L}_{s}$, since the tailoring variables typically constitute a smaller subset of $\overline{L}_{s}$. In addition, (3) suggests an approach to consider both the longitudinal joint score and the full confounder by setting $\overline{b}_{s}=\{\overline{L}_{s},\overline{e}_{s}\}$—the augmented confounder approach. As pointed out in Zigler et al., (2013) and Li et al., (2023), this specification implicitly conducts outcome regression within each propensity score strata and has been considered as a Bayesian analog of a doubly robust estimator (even though it does not strictly possess the frequentist double robustness asymptotic property). In the time-fixed treatment setting, this specification has been shown to improve the efficiency for estimating causal effects compared to adjusting for confounders alone (Hu et al.,, 2021). Finally, Theorem 3 reduces to the g-formula for uncensored longitudinal observational data with deterministic static strategy in Achy-Brou et al., (2010) as a special case when we set $\overline{b}_{s}$ to the history of longitudinal propensity scores.
@@ 4.2 Model specification via Bayesian Additive Regression Trees
To increase the flexibility and robustness of the estimation procedure, we propose to specify and infer the component models of the new g-formula using BART (Chipman et al.,, 2010). Hill, (2011) first introduced BART as a flexible Bayesian nonparametric procedure for estimating the average causal effect with a time-fixed treatment. Several subsequent simulation studies have shown that, compared to other machine learners, causal estimators derived from BART usually have the best performance metrics under a time-fixed treatment setting (see, for example, Hu et al., (2020) for binary outcomes and Hu et al., (2021) for survival outcomes, Dorie et al., (2019) for comparative results from causal inference data analysis challenge launched in the 2016 Atlantic Causal Inference Conference, and Caron et al., (2022) for a recent overview and comparison of methods to estimate conditional average causal effects). In addition, among several advancements in Bayesian nonparametric methods, e.g., the Dirichlet process prior (Antonelli and Daniels,, 2019), the Gaussian process prior (Zhu et al.,, 2023), and the Gamma process prior (Oganisian et al.,, 2024), BART has the additional advantage of computational efficiency with minimum tuning. For the concreteness of method development, we, therefore, describe formal procedures to integrate BART into addressing causal effects with a time-varying treatment effect but acknowledge that other Bayesian nonparametric priors can be similarly integrated into the proposed survival g-formula.
@@ 5.1 Simulation studies with synthetic data
Additional details, including model specifications of the BART estimators and descriptions of other compared estimators, are given in Web Appendix C. We omit the descriptions of the TMLE estimator for brevity and refer readers to Petersen et al., (2014, §4.2.2) for a comprehensive introduction. The NICE, IPW, and AIPW estimators are implemented using R functions available from Wen et al., (2021) and Wen et al., (2022), and the TMLE is implemented via the R package ltmle (Lendle et al.,, 2017) using their default specification. For BART, we draw $15,000$ samples from each posterior distribution, and the first $10,000$ was discarded as burn-in. For each posterior sample of the model parameters, we further generated $10,000$ Monte Carlo samples for the Monte Carlo integration to obtain samples of potential outcomes from the PPD. We replicate the results over 100 synthetic datasets, and values of BART hyper-parameters are set as default, described in Section 4.2. Simulation results in terms of the relative bias (rBIAS) and root mean squared error (RMSE) across five study periods are summarized in Table 1.
@@
Several observations emerge from Table 1. First, the NICE estimator frequently has the largest bias, suggesting that the usual implementation of this estimator is likely to suffer from model misspecification. Second, the IPW, AIPW, and TMLE estimators exhibit similar patterns, as they show minimum biases in some or all of the first three periods but could be considerably biased with inflated RMSEs in the last two periods. Among these three estimators, the IPW gives the largest bias and is the least efficient, but it can be improved, as expected, by the AIPW and TMLE with additional information from outcome modeling. Third, across both levels of censoring, although the proposed BART estimators have slightly larger biases and RMSEs in the first two periods, they have mostly smaller biases and are more efficient in later periods. This is likely because flexible nonparametric generative modeling of balancing scores and outcomes becomes increasingly advantageous as cumulative information on outcomes and covariates becomes available in later periods. Finally, we find that the three BART estimators with different choices of balancing scores generally perform similarly with only slight differences. That is, BART-BS tends to perform better in earlier periods due to its parsimonious modeling structure. In contrast, BART-Cov and BART-Cov-BS exhibit smaller biases and lower RMSE in the final period as the availability of streaming information increases. Moreover, BART-Cov-BS appears to enjoy a slight bias and efficiency advantage over the BART-Cov by conducting outcome regression within each longitudinal joint score strata. This observation resembles the finding in a previous simulation with a time-fixed treatment (Hu et al.,, 2021).
@@ 6 Discussion
Third, the proposed BART estimators, although flexible, are still grounded under the g-formula framework. Hence, by construction, our estimators differ from the double machine learning (DML) or TMLE methods that are grounded under the efficient influence function framework (Petersen et al.,, 2014; Chernozhukov et al.,, 2018). With a time-fixed treatment, Naimi et al., (2021) have shown via simulations that singly robust g-formula estimators with machine learning methods (random forests, extreme gradient boosting, generalized additive models) may not improve the estimation of causal effects over misspecified parametric models, and both are improved by DML or TMLE. In our more complex time-varying treatment scenario, we find that BART estimators can still improve upon misspecified parametric estimators, although they only become less biased and more efficient than AIPW and TMLE at later periods as the availability of streaming information increases. This, along with prior simulations that recommended BART over other machine learning methods (Dorie et al.,, 2019; Hu et al.,, 2020; Caron et al.,, 2022), may suggest that BART remains a competitive data-adaptive approach for causal inference. However, as pointed out by Hill et al., (2020), there are open questions to solve regarding the asymptotic theory of BART. A promising future direction is incorporating BART into the efficient influence function-based estimators, such as DML and TMLE (Antonelli et al.,, 2022); this approach may combine the strengths of BART and rate doubly or multiply robust methods, potentially leading to further improvements over the current methods.

### v4

@@ 1 Introduction
Although the original form of the g-formula is nonparametric, its parametric implementation, based on the maximum likelihood estimates of its component models, has been more commonly used to ensure feasible computation. For instance, by employing generalized linear models, the parametric g-formula estimator has been implemented to adjust for a relatively large number of confounders with either non-censored (Westreich et al.,, 2012) or censored outcomes (Keil et al.,, 2014). However, the parametric g-formula estimator faces two potential challenges. First, for studies involving a moderate number of time-varying confounders, modeling the entire history of time-varying confounders can be a daunting task, especially when the time-varying confounders involve a mix of continuous and categorical variables (Achy-Brou et al.,, 2010). Second, the consistency of the parametric g-formula estimator requires the correct specification of all component models, whereas the need for modeling multi-dimensional confounders in each period is susceptible to model misspecification (Daniel et al.,, 2013). For example, overly parsimonious models can lead to the g-null paradox (Robins and Wasserman,, 1997) and introduce bias in the final causal effect estimates; see McGrath et al., (2022) for further discussion. To this end, methods that improve the parametric g-formula with attention to more flexible model specifications are of increasing interest.
@@
Our second contribution is to integrate the Bayesian Additive Regression Trees (BART), a flexible machine learner, into the survival g-formula to improve upon the standard implementation of the survival parametric g-formula estimator in longitudinal observational studies. Specifically, we incorporate BART for component models in a g-formula to mitigate the bias arising from parametric model misspecification (Chipman et al.,, 2010; Sparapani et al.,, 2016; Tan and Roy,, 2019). We consider this approach because causal estimators based on BART have been empirically demonstrated to outperform competitors in drawing causal inference about the average treatment effect (Hill,, 2011), conditional average treatment effect (Henderson et al.,, 2018; Dorie et al.,, 2019; Hu et al.,, 2021; Hahn et al.,, 2020; Caron et al.,, 2022), as well as causal effects with an intermediate variable (Josefsson and Daniels,, 2021; Bargagli-Stoffi et al.,, 2022; Linero and Zhang,, 2022; Chen et al.,, 2024). This BART approach is flexible as it also provides full posterior samples for all unknown parameters, with built-in uncertainty quantification (Woody et al.,, 2021).
@@ 2.2 Treatment strategies and causal estimands
Finally, we have primarily defined the super-population counterfactual risk estimand within a frequentist framework. Section 2.3 reviews the classical g-formula, which serves to identify this estimand. While frequentist implementations of the g-formula could, in principle, target the super-population estimand, a Bayesian implementation typically approximates it by targeting a mixed counterfactual risk estimand, as discussed in Section 4.1 (see also Li et al., (2023) for a related discussion in the context of a point treatment). In Section 5.1, we evaluate both frequentist and Bayesian g-formula methods under a frequentist framework, assuming the existence of a super-population estimand to ensure a common inferential target.
@@ 2.3 Assumptions and the g-formula—a brief review
Let $f(C_{t+1}=0|\overline{l}_{t},\overline{a}_{t}^{g},C_{t}=Y_{t}=0)$ denote the conditional probability of no censoring in period $t+1$, and then the joint density of treatment status and no censoring, $f(a_{t}^{g},C_{t+1}=0|\overline{l}_{t},\overline{a}_{t-1}^{g},C_{t}=Y_{t}=0)=f(a_{t}^{g}|\overline{l}_{t},\overline{a}_{t-1}^{g},C_{t}=Y_{t}=0)\times f(C_{t+1}=0|\overline{l}_{t},\overline{a}_{t}^{g},C_{t}=Y_{t}=0)>0$, whenever $f(\overline{L}_{t}=\overline{l}_{t},\overline{A}_{t}=\overline{a}_{t}^{g},C_{t}=Y_{t}=0)>0,\forall~{}t\in\mathscr{T}$.
@@ 3 A general class of survival g-formulas
Different specifications of the longitudinal balancing scores can potentially serve different purposes. For example, one can choose the longitudinal joint score or any bijective mapping of these scores as the longitudinal balancing score in (3). This approach could be preferable in some cases because $e_{s}$ is a scalar summary of the higher-dimensional $L_{s}$ in each period $s$, and it is substantially easier to model and then integrate over the one-dimensional longitudinal joint score than multi-dimensional confounders. For a dynamic treatment strategy, one could set $\overline{b}_{s}=\{\overline{H}_{s},\overline{e}_{s}\}$ and maintain a lower dimensionality relative to $\overline{L}_{s}$, since the tailoring variables typically constitute a smaller subset of $\overline{L}_{s}$. In addition, (3) suggests an approach to consider both the longitudinal joint score and the full confounder by setting $\overline{b}_{s}=\{\overline{L}_{s},\overline{e}_{s}\}$—the augmented confounder approach. As pointed out in Zigler et al., (2013) and Li et al., (2023), this specification implicitly conducts outcome regression within each propensity score strata and has been referred to as a “Bayesian analog of a doubly robust estimator” (even though it does not strictly possess the frequentist double robustness asymptotic property). In the time-fixed treatment setting, this specification has been shown to improve the efficiency for estimating causal effects compared to adjusting for confounders alone (Hu et al.,, 2021). Finally, Theorem 3 reduces to the g-formula for uncensored longitudinal observational data with deterministic static strategy in Achy-Brou et al., (2010) as a special case when we set $\overline{b}_{s}$ to the history of longitudinal propensity scores.
@@ 4.2 Model specification via Bayesian Additive Regression Trees
To increase the flexibility and robustness of the estimation procedure, we propose to specify and infer the component models of the new g-formula using BART (Chipman et al.,, 2010). Hill, (2011) first introduced BART as a flexible Bayesian nonparametric procedure for estimating the average causal effect with a time-fixed treatment. Several subsequent simulation studies have shown that, compared to other machine learners, causal estimators derived from BART usually have the best performance metrics under a time-fixed treatment setting (see, for example, Hu et al., (2020) for binary outcomes and Hu et al., (2021) for survival outcomes, Dorie et al., (2019) for comparative results from causal inference data analysis challenge launched in the 2016 Atlantic Causal Inference Conference, and Caron et al., (2022) for a recent overview and comparison of methods to estimate conditional average causal effects). In addition, among several advancements in Bayesian nonparametric methods, e.g., the Dirichlet process prior (Antonelli and Daniels,, 2019), the Gaussian process prior (Zhu et al.,, 2023), and the Gamma process prior (Oganisian et al.,, 2024), BART has the additional advantage of computational efficiency with minimum tuning. For the concreteness of method development, we, therefore, describe procedures to integrate BART into addressing causal effects with a time-varying treatment effect but acknowledge that other Bayesian nonparametric priors can be similarly integrated into the proposed survival g-formula.
@@ 5.1 Simulation studies with synthetic data
Additional details, including model specifications of the BART estimators and descriptions of other compared estimators, are given in Web Appendix C. We omit the descriptions of the TMLE estimator for brevity and refer readers to Petersen et al., (2014, §4.2.2) for a comprehensive introduction. The NICE, IPW, and AIPW estimators are implemented using R functions available from Wen et al., (2021) and Wen et al., (2022), and the TMLE is implemented via the R package ltmle (Lendle et al.,, 2017) using their default specification. As detailed in Web Appendix C, the parametric models (to estimate nuisance functions) underlying the NICE, IPW, and AIPW estimators incorporate only linear covariate adjustments. Consequently, they do not fully capture the nonlinear terms and interactions present in the data-generating model. In contrast, TMLE is a frequentist machine learning estimator that uses flexible data-adaptive methods to represent the nuisance functions and is expected to improve upon parametric implementation of NICE, IPW, and AIPW. This design choice serves to illustrate the comparative performance of common estimators under parametric model misspecification. Similar to TMLE, BART also uses flexible models to represent nuisance functions. To implement BART, we draw $15,000$ samples from each posterior distribution, and the first $10,000$ was discarded as burn-in. For each posterior sample of the model parameters, we further generated $10,000$ Monte Carlo samples for the Monte Carlo integration to obtain samples of potential outcomes from the PPD. We replicate the results over 100 synthetic datasets, and values of BART hyper-parameters are set as default, described in Section 4.2. Simulation results in terms of the relative bias (rBIAS) and root mean squared error (RMSE) across five study periods are summarized in Table 1. To aid in the interpretation of the metric, Web Table 1 in Web Appendix C.7 reports the mean squared error (MSE) along with the Monte Carlo standard error for each MSE estimator, as defined in Morris et al., (2019, Table 6).
@@
Several observations emerge from Table 1. First, the NICE estimator frequently has the largest bias, suggesting that the usual implementation of this estimator is likely to suffer from model misspecification. Second, the IPW, AIPW, and TMLE estimators exhibit similar patterns, as they show minimum biases in some or all of the first three periods but could be considerably biased with inflated RMSEs in the last two periods. Among these three estimators, the IPW gives the largest bias and is the least efficient, but it can be improved, as expected, by the AIPW and TMLE with additional information from outcome modeling. Third, across both levels of censoring, although the proposed BART estimators have slightly larger biases and RMSEs in the first two periods, they have mostly smaller biases and are more efficient in later periods. This is likely because flexible nonparametric generative modeling of balancing scores and outcomes becomes increasingly advantageous as cumulative information on outcomes and covariates becomes available in later periods. Finally, we find that the three BART estimators with different choices of balancing scores generally perform similarly with only slight differences. That is, BART-BS tends to perform better in earlier periods due to its parsimonious modeling structure. In contrast, BART-Cov and BART-Cov-BS exhibit smaller biases and lower RMSE in the final period as the availability of streaming information increases. Moreover, BART-Cov-BS appears to enjoy a slight bias and efficiency advantage over the BART-Cov by conducting outcome regression within each longitudinal joint score strata. This observation resembles the finding in a previous simulation with a time-fixed treatment (Hu et al.,, 2021). However, as shown in Web Table 1, the differences in MSE across different BART implementations sometimes fall within the uncertainty of the simulation experiment.
@@ 6 Discussion
Third, while allowing for flexible estimation of nuisance parameters, when grounded in a frequentist statistical model, the proposed BART estimators do not solve an efficient influence function or share the corresponding properties of double machine learning (DML) or TMLE methods (Petersen et al.,, 2014; Chernozhukov et al.,, 2018). With a time-fixed treatment, Naimi et al., (2021) have shown via simulations that singly robust g-formula estimators with machine learning methods (random forests, extreme gradient boosting, generalized additive models) may not improve the estimation of causal effects over misspecified parametric models, and both are improved by DML or TMLE. In our more complex time-varying treatment scenario, we find that BART estimators can still improve upon misspecified parametric estimators, although they only become less biased and more efficient than AIPW and TMLE at later periods as the availability of streaming information increases. This, along with prior simulations that recommended BART over other machine learning methods (Dorie et al.,, 2019; Hu et al.,, 2020; Caron et al.,, 2022), may suggest that BART remains a competitive data-adaptive approach for causal inference. However, as pointed out by Hill et al., (2020), there are open questions to solve regarding the asymptotic theory of BART, and hence the frequentist inferential properties of BART causal estimators require further investigation. A promising future direction is to incorporate BART as a data-adaptive method to learn the nuisance functions of the efficient influence function tailored to the target causal estimand. Assuming a point treatment, Antonelli et al., (2022) developed an estimator that combines the strengths of frequentist and Bayesian methods for causal inference and proposed a valid confidence interval estimator. Extending their estimator to accommodate time-varying treatments could lead to an estimator that combines the advantages of BART with rate-doubly or multiply robust methods, offering improvements over the current BART-based g-formula estimators.